Suppose your firm earns $9 million in taxable income.

What is the firm's tax liability?

What is the average tax rate?

What is the marginal tax rate?

50,000 15%

50,001 - 75,000 25%

75,001- 100,000. 34%

100,001-335,000. 39%

335,001-10,000,000. 34%

10,000,001-15,000,000. 35%

15,000,001-18,333,333. 38%

18,333,334- 34%


1
Expert's answer
2021-04-14T14:47:08-0400

a) Firms tax liability

= ( 50,000* 0.15)+(25,000* 0.25)+(25,000* 0.34)+(235,000* 0.39)+(8,665,000* 0.34)

= 7,500+6,250+8,500+91,650+2,946,100

= 3,060,000

b) Average tax rate

= 3,060,000/9,000,000

= 0.34

= 34%

c)Marginal tax rate

= 34%